Find the value of the constant c , if the difference of the roots of x^2-11x+c = 0
is 3

Answer by vksarvepalli(154)

x^2-11x+c = 0 let a,b be the roots of this equation.
sum of the roots is (a+b)=11
product of the roots ab = c
now given (a-b)=3
(a-b)^2=(a+b)^2-4ab
-> 9 = 121-4c
=> 4c = 112
=> c=28
ans. c=28